what is the capacity of a spherical gas tank that is 45 feet in diameter? i know the formula (4/3)(3.14)(r)^3, i just dont know how to use it!
Answer by feliz1965(151)

Let C = capacity of spherical gas tank.
You are given a 45 foot diameter. We need to find r, which is the radius.
The radius is 1/2 the diameter.
So, radius is 45/2 = 22.5
C = (4/3)(3.14)(22.5)^3
Before finding C (your answer), we must raise 22.5 to the third power.
22.5 x 22.5 x 22.5 = 11390.625
Now, C = (4/3)(3.14)(11390.625)
Of course, we can also write 3/4 as 0.75 giving us the equation
C = (0.75)(3.14)(11390.625)
Can you finish now?
Just multiply using your calculator to find C.